# NoRA: Breaking the Linear Ceiling of Low-Rank Adaptation via Manifold Expansion

### v1

@@ 5.2.2 Activation: The Necessity of Non-linearity
Row (b) validates the “non-linear” in NoRA. Replacing SiLU with an identity function (effectively reverting to a linear parallel adapter) yields the worst PPL (3.97). Furthermore, SiLU outperforms ReLU (3.84), suggesting that the smooth gating property of SiLU facilitates better gradient flow and manifold curvature than the rigid sparsity of ReLU.
@@ 5.2.3 Dropout: Structural Manifold Expansion
Row (c) reveals a critical insight: removing dropout ($p=0$) hurts generalization (PPL increases from 3.81 to 3.85). This confirms that in NoRA, dropout acts as a structural manifold expander. By stochastically blocking latent paths, it forces the model to distribute information across more dimensions, preventing the network from over-relying on dominant singular values.
@@ 5.3 Efficiency Trade-off: Capacity per Parameter
Table 3: Efficiency-Performance Trade-off: Comparison on SlimOrca. (1) Parameter efficiency: NoRA ($r=64$) matches the performance of LoRA ($r=512$) with $\approx 8\times$ fewer parameters. (2) Optimal Balance: NoRA ($r=128$) achieves a great perplexity (3.81) while using only 25% of the LoRA baseline’s parameters. (3) Latency: Inference overhead is dominated by kernel launching rather than computation size, resulting in consistent throughput ($\approx 51$ tok/s) across NoRA ranks.
@@
NoRA 64 27.3M 3.89 $1.06\times$ 50.90 tok/s
NoRA 128 54.5M 3.81 $1.06\times$ 51.05 tok/s
@@ 5.3.1 Parameter Efficiency
Table 3 reveals a compelling contrast in parameter utilization. NoRA at rank 64 achieves a perplexity of 3.89, effectively matching the LoRA baseline at rank 512 (3.90) while using only 27.3M parameters, an $8\times$ reduction compared to LoRA’s 218.1M. Furthermore, increasing the NoRA rank to 128 (54.5M parameters) reduces the perplexity to 3.81, significantly outperforming the high-rank LoRA baseline while still consuming only 25% of its parameter budget. This confirms that structural complexity (non-linearity) is a significantly more efficient driver of performance than brute-force dimensional scale.
@@ 5.3.2 Inference Latency
Since NoRA involves non-linear operations, the weights cannot be mathematically merged into the base model ($W_{0}+\Delta W$) for zero-latency inference. However, in modern multi-tenant serving systems (e.g., S-LoRA, Punica), unmerged inference is the standard paradigm to support dynamic adapter switching.
Our benchmarks indicate a modest latency overhead of approximately 6% compared to the merged LoRA baseline. Notably, the throughput for NoRA remains consistent across ranks 64 and 128 ($\approx 51$ tok/s). This suggests that the overhead is primarily driven by fixed kernel launching costs (for SiLU and Dropout operations) rather than the computational cost of the adapter’s matrix multiplication. We argue that for high-stakes reasoning tasks, this marginal latency cost is well justified by the significant gains in generation quality and parameter efficiency.
@@ 6 Related Work
We position NoRA within the PEFT landscape by contrasting it with linear low-rank variants and existing adapter architectures. Table 4 summarizes the structural differences.
Table 4: Feature Comparison. Unlike LoRA and its variants, NoRA introduces nonlinearity. Unlike traditional parallel adapters, which operate at the coarse module-level, NoRA operates at the fine-grained weight-level, enabling precise manifold deformation within attention projections.
Feature LoRA Houlsby Adapter Parallel Adapter NoRA (Ours)
@@ 6.1 LoRA and Linear Variants
LoRA [5] established the standard for efficient fine-tuning by hypothesizing that weight updates reside in a low-rank linear subspace. Subsequent work has optimized this paradigm: AdaLoRA [11] allocates rank budgets dynamically; DoRA [7] decomposes updates into magnitude and direction for stability; and QLoRA [2] focuses on quantization. Crucially, while these methods improve efficiency or stability, they remain bound by the linear form $\Delta W=BA$. Our work identifies this linearity as the primary bottleneck for reasoning-intensive tasks. NoRA is orthogonal to these optimizations; it challenges the functional form of the update itself, replacing the linear subspace assumption with a non-linear manifold hypothesis.
@@ 6.2 Granularity in Parallel Adapters
The original series of adapters [4] inserted MLPs sequentially, resulting in high latency. To mitigate this, parallel adapters [3, 12] placed adapters along transformer sub-layers. However, a critical distinction lies in granularity. Prior parallel adapters operate at the module-level – processing the aggregate output of an entire attention or feedforward network (FFN) block (visualized as the “Parallel Adapter” in Figure 1). In contrast, NoRA operates at the weight level, injecting nonlinearity directly into the internal query ($W_{q}$) and value ($W_{v}$) projections. This fine-grained intervention allows NoRA to fundamentally alter the attention mechanism’s internal feature dynamics rather than merely correcting its output.
@@ 6.3 The Expressivity-Mergeability Trade-off
In the PEFT literature, linearity is often preserved solely to allow weight merging ($\Delta W+W$) for zero-latency inference. We argue that for high-value vertical domains (e.g., mathematics, logic), the expressivity gains from nonlinearity outweigh the cost of unmerged inference. Furthermore, with the rise of multi-tenant serving systems such as S-LoRA [9] and Punica [1], which dynamically serve unmerged adapters, the deployment barrier for non-linear adapters such as NoRA has been effectively removed.
@@
Recent advances in PEFT have focused on optimizing the learning dynamics of low-rank updates. In particular, DoRA [7] decomposes weights into magnitude and direction components, successfully bridging the gap between LoRA and full fine-tuning. However, DoRA and similar variants (e.g., AdaLoRA [11]) fundamentally operate within the linear adaptation paradigm, where the final effective weight remains $W^{\prime}=W+\Delta W$. In contrast, NoRA pursues an orthogonal direction: structural expansion. By introducing non-linear gating (SiLU), dropout, and MLP-like interactions, NoRA breaks the linear ceiling, expanding the functional class of the adapter itself. While DoRA optimizes how we learn a linear subspace, NoRA changes what the subspace can represent.
@@ 7 Limitations
Although NoRA demonstrates significant advantages in reasoning-intensive domains, we acknowledge certain limitations. First, the requirement for unmerged inference may pose challenges for edge devices with strictly limited memory bandwidth, where weight merging is mandatory. Second, our current analysis focuses on the Llama-3 architecture; future work is needed to validate the universality of the “linear ceiling” hypothesis across other architectures like Mixture-of-Experts (MoE) or Mamba.
@@ 8 Conclusion
We introduce NoRA (Non-linear Rank Adaptation), a weight-level architecture that injects SiLU gating and structural dropout directly into the low-rank bottleneck. Our findings demonstrate that with sufficient data, such as in the SlimOrca benchmark, NoRA unlocks the latent capacity that linear methods waste. Crucially, NoRA at rank 64 outperforms LoRA at rank 512, demonstrating that non-linearity enables superior spectral efficiency. Through singular value spectrum analysis, we further confirm that NoRA activates the dormant tail of the singular value spectrum, significantly increasing the effective rank and preventing the representation collapse observed in linear updates.
We argue that for high-value vertical domains, the rigorous demands of reasoning outweigh the convenience of weight merging. NoRA proves that non-linearity is key to unlocking high-rank potential, positioning itself not merely as another adapter variant, but as a necessary evolution for scenarios where reasoning depth is paramount. Future work will extend this non-linear paradigm to larger-scale backbones (e.g., 70B parameters) and multi-modal reasoning tasks.
Finally, we note that the structural benefits of NoRA are theoretically orthogonal to the parameterization improvements of DoRA. A promising direction for future work is to explore a hybrid “Weight-Decomposed Non-linear Adapter,” potentially combining the stable optimization dynamics of DoRA with the high expressivity of NoRA.
@@ Appendix A Extended Spectral Analysis on SlimOrca
Figure 6: Cumulative Spectral Energy (AUC-90) on SlimOrca. LoRA reaches energy saturation rapidly, indicating that the majority of the rank budget is redundant. In contrast, NoRA’s manifold dimensionality continues growing, demonstrating a more uniform distribution of information across singular dimensions.
In the main text, we demonstrate NoRA’s performance superiority on the SlimOrca benchmark. Here, we investigate whether the theoretical mechanism – manifold expansion – holds for this general reasoning dataset.
@@
  • NoRA’s Expansion: Conversely, NoRA continues growing as rank increases. This means that the spectral energy is distributed across a broader range of singular vectors. The non-linear gating mechanism effectively forces the optimization process to utilize the “tail” dimensions, thereby expanding the representation manifold to capture the diverse reasoning patterns found in SlimOrca.
@@ Appendix B Comprehensive Spectral Evaluation on MathInstruct
Mathematical reasoning requires high-precision state tracking, making it an ideal testbed for analyzing the expressivity of adapters. We present a detailed spectral breakdown of NoRA versus LoRA on the MathInstruct dataset.
@@ B.1 Singular Value Spectrum Analysis (SVSA)
Figure 7: Spectral Signature of the MathInstruct dataset. LoRA exhibits “rank collapse,” where singular values drop precipitously, indicating under-utilization of the rank budget. NoRA maintains a “heavy tail,” activating a broader subspace.
@@
  • Manifold Expansion in NoRA: NoRA maintains a significantly higher magnitude in the tail of the spectrum (the “heavy tail”). This spectral signature proves that the non-linear components (SiLU, Dropout, and MLP structures) prevent the weight updates from collapsing into a trivial low-rank subspace, thereby breaking the linear ceiling.
@@ B.2 Effective Rank and Information Capacity
Figure 8: Effective Rank Analysis on MathInstruct. We compare the spectral utilization of LoRA and NoRA across increasing rank budgets ($r\in\{16,64,128,512\}$). As the rank budget increases to 512, LoRA’s effective rank saturates significantly, whereas NoRA’s effective rank scales with the budget.
@@
To quantify the “richness” of the learned representation, we compute the effective rank ($e^{H}$, where $H$ is the spectral entropy). As shown in Figure 8, NoRA achieves a substantially higher effective rank compared to LoRA configurations of similar or even larger parameter counts. This metric serves as a proxy for information capacity, suggesting that NoRA encodes more complex functional transformations within the same dimensional budget.
@@ B.3 Spectral Energy Distribution
Figure 9: AUC-90 on MathInstruct. NoRA requires more singular components to reach 90% energy variance, confirming a richer and less sparse representation compared to LoRA.
@@
Figure 9 corroborates the findings of the SVSA and the effective rank analysis. Similarly to the observation in SlimOrca, the LoRA baseline on MathInstruct suffers from rapid energy saturation. NoRA’s manifold dimensionality grows as rank increases, indicating that the model relies on a diverse set of orthogonal features to solve mathematical problems, rather than overfitting to a few dominant directions.

### v2

@@ 5.2.2 Activation: The Necessity of Non-linearity
Row (b) validates the “non-linear” in CeRA. Replacing SiLU with an identity function (effectively reverting to a linear parallel adapter) yields the worst PPL (3.97). Furthermore, SiLU outperforms ReLU (3.84), suggesting that the smooth gating property of SiLU facilitates better gradient flow and manifold curvature than the rigid sparsity of ReLU.
@@ 5.2.3 Dropout: Structural Manifold Expansion
Row (c) reveals a critical insight: removing dropout ($p=0$) hurts generalization (PPL increases from 3.81 to 3.85). This confirms that in CeRA, dropout acts as a structural manifold expander. By stochastically blocking latent paths, it forces the model to distribute information across more dimensions, preventing the network from over-relying on dominant singular values.
@@ 5.3 Efficiency Trade-off: Capacity per Parameter
Table 3: Efficiency-Performance Trade-off: Comparison on SlimOrca. (1) Parameter efficiency: CeRA ($r=64$) matches the performance of LoRA ($r=512$) with $\approx 8\times$ fewer parameters. (2) Optimal Balance: CeRA ($r=128$) achieves a great perplexity (3.81) while using only 25% of the LoRA baseline’s parameters. (3) Latency: Inference overhead is dominated by kernel launching rather than computation size, resulting in consistent throughput ($\approx 51$ tok/s) across CeRA ranks.
@@
CeRA 64 27.3M 3.89 $1.06\times$ 50.90 tok/s
CeRA 128 54.5M 3.81 $1.06\times$ 51.05 tok/s
@@ 5.3.1 Parameter Efficiency
Table 3 reveals a compelling contrast in parameter utilization. CeRA at rank 64 achieves a perplexity of 3.89, effectively matching the LoRA baseline at rank 512 (3.90) while using only 27.3M parameters, an $8\times$ reduction compared to LoRA’s 218.1M. Furthermore, increasing the CeRA rank to 128 (54.5M parameters) reduces the perplexity to 3.81, significantly outperforming the high-rank LoRA baseline while still consuming only 25% of its parameter budget. This confirms that structural complexity (non-linearity) is a significantly more efficient driver of performance than brute-force dimensional scale.
@@ 5.3.2 Inference Latency
Since CeRA involves non-linear operations, the weights cannot be mathematically merged into the base model ($W_{0}+\Delta W$) for zero-latency inference. However, in modern multi-tenant serving systems (e.g., S-LoRA, Punica), unmerged inference is the standard paradigm to support dynamic adapter switching.
Our benchmarks indicate a modest latency overhead of approximately 6% compared to the merged LoRA baseline. Notably, the throughput for CeRA remains consistent across ranks 64 and 128 ($\approx 51$ tok/s). This suggests that the overhead is primarily driven by fixed kernel launching costs (for SiLU and Dropout operations) rather than the computational cost of the adapter’s matrix multiplication. We argue that for high-stakes reasoning tasks, this marginal latency cost is well justified by the significant gains in generation quality and parameter efficiency.
@@ 6 Related Work
We position CeRA within the PEFT landscape by contrasting it with linear low-rank variants and existing adapter architectures. Table 4 summarizes the structural differences.
Table 4: Feature Comparison. Unlike LoRA and its variants, CeRA introduces nonlinearity. Unlike traditional parallel adapters, which operate at the coarse module-level, CeRA operates at the fine-grained weight-level, enabling precise manifold deformation within attention projections.
Feature LoRA Houlsby Adapter Parallel Adapter CeRA (Ours)
@@ 6.1 LoRA and Linear Variants
LoRA [5] established the standard for efficient fine-tuning by hypothesizing that weight updates reside in a low-rank linear subspace. Subsequent work has optimized this paradigm: AdaLoRA [11] allocates rank budgets dynamically; DoRA [7] decomposes updates into magnitude and direction for stability; and QLoRA [2] focuses on quantization. Crucially, while these methods improve efficiency or stability, they remain bound by the linear form $\Delta W=BA$. Our work identifies this linearity as the primary bottleneck for reasoning-intensive tasks. CeRA is orthogonal to these optimizations; it challenges the functional form of the update itself, replacing the linear subspace assumption with a non-linear manifold hypothesis.
@@ 6.2 Granularity in Parallel Adapters
The original series of adapters [4] inserted MLPs sequentially, resulting in high latency. To mitigate this, parallel adapters [3, 12] placed adapters along transformer sub-layers. However, a critical distinction lies in granularity. Prior parallel adapters operate at the module-level – processing the aggregate output of an entire attention or feedforward network (FFN) block (visualized as the “Parallel Adapter” in Figure 1). In contrast, CeRA operates at the weight level, injecting nonlinearity directly into the internal query ($W_{q}$) and value ($W_{v}$) projections. This fine-grained intervention allows CeRA to fundamentally alter the attention mechanism’s internal feature dynamics rather than merely correcting its output.
@@ 6.3 The Expressivity-Mergeability Trade-off
In the PEFT literature, linearity is often preserved solely to allow weight merging ($\Delta W+W$) for zero-latency inference. We argue that for high-value vertical domains (e.g., mathematics, logic), the expressivity gains from nonlinearity outweigh the cost of unmerged inference. Furthermore, with the rise of multi-tenant serving systems such as S-LoRA [9] and Punica [1], which dynamically serve unmerged adapters, the deployment barrier for non-linear adapters such as CeRA has been effectively removed.
@@
Recent advances in PEFT have focused on optimizing the learning dynamics of low-rank updates. In particular, DoRA [7] decomposes weights into magnitude and direction components, successfully bridging the gap between LoRA and full fine-tuning. However, DoRA and similar variants (e.g., AdaLoRA [11]) fundamentally operate within the linear adaptation paradigm, where the final effective weight remains $W^{\prime}=W+\Delta W$. In contrast, CeRA pursues an orthogonal direction: structural expansion. By introducing non-linear gating (SiLU), dropout, and MLP-like interactions, CeRA breaks the linear ceiling, expanding the functional class of the adapter itself. While DoRA optimizes how we learn a linear subspace, CeRA changes what the subspace can represent.
@@ 7 Limitations
Although CeRA demonstrates significant advantages in reasoning-intensive domains, we acknowledge certain limitations. First, the requirement for unmerged inference may pose challenges for edge devices with strictly limited memory bandwidth, where weight merging is mandatory. Second, our current analysis focuses on the Llama-3 architecture; future work is needed to validate the universality of the “linear ceiling” hypothesis across other architectures like Mixture-of-Experts (MoE) or Mamba.
@@ 8 Conclusion
We introduce CeRA (Capacity-enhanced Rank Adaptation), a weight-level architecture that injects SiLU gating and structural dropout directly into the low-rank bottleneck. Our findings demonstrate that with sufficient data, such as in the SlimOrca benchmark, CeRA unlocks the latent capacity that linear methods waste. Crucially, CeRA at rank 64 outperforms LoRA at rank 512, demonstrating that non-linearity enables superior spectral efficiency. Through singular value spectrum analysis, we further confirm that CeRA activates the dormant tail of the singular value spectrum, significantly increasing the effective rank and preventing the representation collapse observed in linear updates.
We argue that for high-value vertical domains, the rigorous demands of reasoning outweigh the convenience of weight merging. CeRA proves that non-linearity is key to unlocking high-rank potential, positioning itself not merely as another adapter variant, but as a necessary evolution for scenarios where reasoning depth is paramount. Future work will extend this non-linear paradigm to larger-scale backbones (e.g., 70B parameters) and multi-modal reasoning tasks.
Finally, we note that the structural benefits of CeRA are theoretically orthogonal to the parameterization improvements of DoRA. A promising direction for future work is to explore a hybrid “Weight-Decomposed Non-linear Adapter,” potentially combining the stable optimization dynamics of DoRA with the high expressivity of CeRA.
@@ Appendix A Extended Spectral Analysis on SlimOrca
Figure 6: Cumulative Spectral Energy (AUC-90) on SlimOrca. LoRA reaches energy saturation rapidly, indicating that the majority of the rank budget is redundant. In contrast, CeRA’s manifold dimensionality continues growing, demonstrating a more uniform distribution of information across singular dimensions.
In the main text, we demonstrate CeRA’s performance superiority on the SlimOrca benchmark. Here, we investigate whether the theoretical mechanism – manifold expansion – holds for this general reasoning dataset.
@@
  • CeRA’s Expansion: Conversely, CeRA continues growing as rank increases. This means that the spectral energy is distributed across a broader range of singular vectors. The non-linear gating mechanism effectively forces the optimization process to utilize the “tail” dimensions, thereby expanding the representation manifold to capture the diverse reasoning patterns found in SlimOrca.
@@ Appendix B Comprehensive Spectral Evaluation on MathInstruct
Mathematical reasoning requires high-precision state tracking, making it an ideal testbed for analyzing the expressivity of adapters. We present a detailed spectral breakdown of CeRA versus LoRA on the MathInstruct dataset.
@@ B.1 Singular Value Spectrum Analysis (SVSA)
Figure 7: Spectral Signature of the MathInstruct dataset. LoRA exhibits “rank collapse,” where singular values drop precipitously, indicating under-utilization of the rank budget. CeRA maintains a “heavy tail,” activating a broader subspace.
@@
  • Manifold Expansion in CeRA: CeRA maintains a significantly higher magnitude in the tail of the spectrum (the “heavy tail”). This spectral signature proves that the non-linear components (SiLU, Dropout, and MLP structures) prevent the weight updates from collapsing into a trivial low-rank subspace, thereby breaking the linear ceiling.
@@ B.2 Effective Rank and Information Capacity
Figure 8: Effective Rank Analysis on MathInstruct. We compare the spectral utilization of LoRA and CeRA across increasing rank budgets ($r\in\{16,64,128,512\}$). As the rank budget increases to 512, LoRA’s effective rank saturates significantly, whereas CeRA’s effective rank scales with the budget.
@@
To quantify the “richness” of the learned representation, we compute the effective rank ($e^{H}$, where $H$ is the spectral entropy). As shown in Figure 8, CeRA achieves a substantially higher effective rank compared to LoRA configurations of similar or even larger parameter counts. This metric serves as a proxy for information capacity, suggesting that CeRA encodes more complex functional transformations within the same dimensional budget.
@@ B.3 Spectral Energy Distribution
Figure 9: AUC-90 on MathInstruct. CeRA requires more singular components to reach 90% energy variance, confirming a richer and less sparse representation compared to LoRA.
@@
Figure 9 corroborates the findings of the SVSA and the effective rank analysis. Similarly to the observation in SlimOrca, the LoRA baseline on MathInstruct suffers from rapid energy saturation. CeRA’s manifold dimensionality grows as rank increases, indicating that the model relies on a diverse set of orthogonal features to solve mathematical problems, rather than overfitting to a few dominant directions.